OpenStudy (anonymous):

Find the length of the hypotenuse of a 45°- 45°-90° triangle with a leg of 8.

OpenStudy (zale101):

use the Pythagorean theorem

OpenStudy (zale101):

a^2+b^2=c^2 Pythagorean theorem

OpenStudy (zale101):

the legs will have equal lengths since you have 45°- 45°-90° triangle

OpenStudy (zale101):

@tylerj37, you're with me?

OpenStudy (anonymous):

ok i think i got 8. 8 or 8x square root of 2

OpenStudy (anonymous):

and yes

OpenStudy (zale101):

8* sqrt(2), there you go (8)^2 + (8)^2 = 64 + 64 = 128 sqrt(128) = 8* sqrt(2)

OpenStudy (anonymous):

awesome thank you!

OpenStudy (zale101):

anytime ^_^

OpenStudy (anonymous):

ok i have another one